# On the Existence and Behavior of Secondary Attention Sinks

### v2

@@ Abstract
Attention sinks are tokens, often the beginning-of-sequence (BOS) token, that receive disproportionately high attention despite limited semantic relevance. In this work, we identify a class of attention sinks, which we term secondary sinks, that differ fundamentally from the sinks studied in prior works, which we term primary sinks. While prior works have identified that tokens other than BOS can sometimes become sinks, they were found to exhibit properties analogous to the BOS token. Specifically, they emerge at the same layer, persist throughout the network and draw a large amount of attention mass. Whereas, we find the existence of secondary sinks that arise primarily in middle layers and can persist for a variable number of layers, and draw a smaller, but still significant, amount of attention mass. Through extensive experiments across 11 model families, we analyze where these secondary sinks appear, their properties, how they are formed, and their impact on the attention mechanism. Specifically, we show that: (1) these sinks are formed by specific middle-layer MLP modules; these MLPs map token representations to vectors that align with the direction of the primary sink of that layer. (2) The $\ell_{2}$-norm of these vectors determines the sink score of the secondary sink, and also the number of layers it lasts for, thereby leading to different impacts on the attention mechanisms accordingly. (3) The primary sink weakens in middle layers, coinciding with the emergence of secondary sinks. We observe that in larger-scale models, the location and lifetime of the sinks, together referred to as sink levels, appear in a more deterministic and frequent manner. Specifically, we identify three sink levels in QwQ-32B and six levels in Qwen3-14B.
@@ 1 Introduction
In this work, we show that all of the multiple tokens that act as attention sinks can be organized into distinct sink levels. The primary level normally corresponds to the BOS sink: it emerges at the same layer as the BOS token and persists throughout the network. Additional sink levels arise in the middle layers and persist for a variable number of layers; we refer to these as Secondary Sinks. Subsequent sections study these secondary sinks in detail.
@@
  • Secondary Sinks shares a similar direction with the primary sink. This direction is encoded in specific middle-layer $l_{\text{start}}$ MLP modules, which convert multiple orthogonal directions to the same sink direction. After $l_{\text{start}}$, a set of semantically uninformative tokens are transformed into attention sinks. Meanwhile, the layers preceding $l_{\text{start}}$ play a key role in constructing this set, distinguishing these tokens from other semantically uninformative tokens.
@@ 2 Background and Preliminary
Each decoder layer $l$ consists of a multi-head self-attention (MHSA) module and a multi-layer perceptron (MLP) , both of which operate on the decoder input hidden state, also referred to as the residual stream ${\bm{H}}^{l}\in{\mathbb{R}}^{t\times h}$. The MHSA produces an output ${\bm{O}}^{l}\in{\mathbb{R}}^{t\times h}$, and the MLP produces ${\bm{F}}^{l}\in{\mathbb{R}}^{t\times h}$; in both cases, the outputs are added back to the residual stream. Decoders may use either a pre-norm or post-norm architecture for which the normalization is applied either before or after each module. The majority of modern models employ pre-norm as shown in Figure˜4.
@@ 3 Properties of Secondary Sinks
Figure˜1 (top) shows the $\ell_{2}$ norms of the hidden states, keys, and values of attention sinks, compared with those of an average token in DeepSeek-14B denoted by $\bar{{\bm{h}}}_{t\neq\text{sink}}$. Token ${\bm{h}}_{0}$ corresponds to the BOS token, while tokens ${\bm{h}}_{1}$ and ${\bm{h}}_{2}$ are tokens from the chat template. We classify all of these sinks as primary sinks, as they emerge simultaneously and persist for the same number of layers as the BOS sink. In addition to these BOS sinks, we observe secondary sinks at positions 1120, 2768, and 9530. These tokens behave like normal tokens up to layer 22, after which they are converted into sinks. Before layer 22, their hidden-state, key, and value norms follow the average-token trend; afterward, they closely track the BOS sink, whose hidden-state norm is several orders of magnitude higher than that of normal tokens, while its key and value norms are lower.
Not all models exhibit secondary sinks. For example, although multiple sinks also appear in LLaMA-7B-Chat Figure˜1 (Bottom), they are exclusively primary sinks. Among the 11 model families we analyzed, secondary sinks were found in Qwen2, Qwen2-Math, Qwen2.5, Qwen2.5-Math, Qwen3, and QwQ. Full detection results on all 11 model families can be found in Appendix A. These sinks typically occur in semantically uninformative tokens Table˜1 and can appear at any position during the generation Figure˜3.
@@
Figure 3: Distribution of sinks across token index.
Model Secondary Sinks Token Distribution
Deepseek-14b " " "1" "^ " "2" "," "(" " I"
56.41% 17.95% 5.13% 5.13% 2.56% 2.56% 2.56%
[…]
19.87% 19.87% 13.91% 9.93% 4.64% 3.97% 3.97%
[…]
6.13% 5.28% 3.62% 3.56% 3.50% 3.48% 2.98%
Qwen2.5-14B " " "0" "1" "8" "=" "_" "t"
74.71% 9.20% 5.75% 2.30% 2.30% 1.15% 1.15%
Table 1: Frequency of secondary sink tokens.
@@
Finding 1.1: Middle Layer Emergence $\triangleright$ Figure˜1 Unlike the typical BOS sink tokens that appear in the early layers of the network, secondary sinks tend to emerge in the middle layers of LLMs. Finding 1.2: Model Size Effect on Sink Levels $\triangleright$ Figure˜2 Larger models tend to exhibit fewer but more consistent and clearly differentiated levels.
@@ 4 Causal Formation Of Secondary Sinks
Figure˜4 (Right) shows the cosine similarity analysis of the 75 future secondary sink tokens collected from AIME24 on DeepSeek-14B at layer 22. The MLP inputs ${\bm{x}}_{t}$ initially exhibit relatively low similarity to the BOS token, but as they pass through the MLP, their cosine similarity gradually increases. By the output of the MLP ${\bm{f}}_{t}$, the vectors are almost fully aligned with the BOS attention direction.
These results indicate that the sink direction is encoded within the MLP layers, and that future secondary sink tokens contain the information needed to recover this direction. Motivated by the linear hypothesis (Park et al., 2023), we hypothesize that these future secondary sinks share a common key can retrieve the sink direction.
To validate this hypothesis, we perform PCA on the concatenated MLP input matrix ${\bm{X}}^{l}$ constructed from 250 sink tokens. We find that the resulting PCA spectrum is highly low-rank. We then feed each pincipal component individually through the MLP and find that multiple different principal components are converted into the same sink direction. Figure˜5 shows the MLP output norm and its cosine similarity with the BOS sink token when feeding $\pm\alpha\mathrm{PC}_{i}$ as input. Here we scale each $\mathrm{PC}_{i}$ by $\pm\alpha$ to stimulate the magnitude of $\mathrm{PC}_{i}$ of each sink. We observe that for each $\mathrm{PC}_{i}$, $\alpha\mathrm{PC}_{i}$ and $\alpha\mathrm{PC}_{i}$ produce MLP outputs that either aligned or misaligned with the BOS sink direction. Crucially, the MLP selectively amplifies the component aligned with the BOS sink, producing the large $\ell_{2}$-norms observed in the hidden states.
@@
Token Swapping
Finding 2.1: Emergence from MLP Layers $\triangleright$ Figures˜4 and 5 The secondary sinks primarily emerge from specific MLP layers in the network, as evidenced by an upsurge in cosine similarity to the BOS sink token after these layers. Finding 2.2: Progressive Formation $\triangleright$ Figures˜6 and 8 The formation of secondary sinks can be traced back to very early layers of the model.
@@ 5 Secondary Sinks Effect on Attention
As shown in  Figure˜7 (Left) , the sink-score for the BOS sink shows a valley shape across the model depth, and it is when it reaches its lowest sink score that the secondary sinks started emerging. This suggests the secondary sinks may serve to compensate for the decay in the BOS sink.
@@
We observe that secondary sinks persist for varying depths, ranging from as few as 2 layers to as many as 22 layers, approximately half of the network. As shown in Figure˜7 (right), both the sink lifetime and the sink score of the secondary sink are strongly correlated with the $\ell_{2}$-norm of the MLP output ${\bm{f}}^{22}_{t}$ at $l_{\text{start}}$. The ratio between the sink scores of the secondary sink and the BOS sink exhibits a log-linear relationship with ${\bm{f}}^{22}_{t}$, with the largest slope observed at layer 33. In contrast, the sink lifetime increases monotonically with the log-norm, exhibiting distinct linear regimes separated by plateaus.
Finding 3.1: Compensating Sink-score $\triangleright$ Figure˜7 (Left) The BOS sink-score exhibits a valley-shaped profile across model depth; as it reaches its minimum, secondary sinks emerge, compensating for the decay of BOS sink strength. Finding 3.2: Sink-score & Lifetime $\triangleright$ Figure˜7 (Mid, Right) The sink-score and lifetime of secondary sinks are both correlated with the norm of MLP outputs. A higher MLP output norm corresponds to a higher sink-score and longer lifetime.

### v3

@@ Abstract
Attention sinks are tokens, often the beginning-of-sequence (BOS) token, that receive disproportionately high attention despite limited semantic relevance. In this work, we identify a class of attention sinks, which we term secondary sinks, that differ fundamentally from the sinks studied in prior works, which we term primary sinks. While prior works have identified that tokens other than BOS can sometimes become sinks, they were found to exhibit properties analogous to the BOS token. Specifically, they emerge at the same layer, persist throughout the network and draw a large amount of attention mass. Whereas, we find the existence of secondary sinks that arise primarily in middle layers and can persist for a variable number of layers, and draw a smaller, but still significant, amount of attention mass. Through extensive experiments across 11 model families, we analyze where these secondary sinks appear, their properties, how they are formed, and their impact on the attention mechanism. Specifically, we show that: (1) these sinks are formed by specific middle-layer MLP modules; these MLPs map token representations to vectors that align with the direction of the primary sink of that layer, (2) the $\ell_{2}$-norm of these vectors determines the sink score of the secondary sink, and also the number of layers it lasts for, thereby leading to different impacts on the attention mechanisms accordingly, (3) the primary sink weakens in the middle layers, coinciding with the emergence of secondary sinks. We observe that in larger-scale models, the location and lifetime of the sinks, together referred to as sink levels, appear in a more deterministic and frequent manner. Specifically, we identify three sink levels in QwQ-32B and six levels in Qwen3-14B. We open-sourced our findings at github.com/JeffreyWong20/Secondary-Attention-Sinks.
@@ 1 Introduction
In this work, we show that all of the multiple tokens that act as attention sinks can be organized into distinct sink levels. The primary level normally corresponds to the BOS sink: it emerges at the same layer as the BOS token and persists throughout the network. Additional sink levels arise in the middle layers and persist for a variable number of layers; we refer to these as Secondary Sinks.
@@
  • Secondary Sinks shares a similar direction with the primary sink. This direction is encoded in specific middle-layer $l_{\text{start}}$ MLP modules, which convert multiple orthogonal directions to the same sink direction. After $l_{\text{start}}$, a set of semantically uninformative tokens are transformed into attention sinks. The layers preceding $l_{\text{start}}$ play a key role in constructing this set, distinguishing these tokens from other semantically uninformative tokens.
@@ 2 Background and Preliminary
Each decoder layer $l$ consists of a multi-head self-attention (MHSA) module and a multi-layer perceptron (MLP) , both of which operate on the decoder input hidden state, also referred to as the residual stream ${\bm{H}}^{l}\in{\mathbb{R}}^{t\times h}$. The MHSA produces an output ${\bm{O}}^{l}\in{\mathbb{R}}^{t\times h}$, and the MLP produces ${\bm{F}}^{l}\in{\mathbb{R}}^{t\times h}$; in both cases, the outputs are added back to the residual stream. Decoders may use either a pre-norm or post-norm architecture for which the normalization is applied either before or after each module. The majority of modern models employ pre-norm as shown in Figure˜3.
@@ 3 Properties of Secondary Sinks
Figure˜1 (top) shows the $\ell_{2}$-norms of the hidden states, keys, and values of attention sinks, compared with those of an average token in DeepSeek-14B denoted by $\bar{{\bm{h}}}_{t\neq\text{sink}}$. Token ${\bm{h}}_{0}$ corresponds to the BOS token, while tokens ${\bm{h}}_{1}$ and ${\bm{h}}_{2}$ are tokens from the chat template. We classify all of these sinks as primary sinks, as they emerge simultaneously and persist for the same number of layers as the BOS sink. In addition to these BOS sinks, we observe secondary sinks at positions 1120, 2768, and 9530. These tokens behave like normal tokens up to layer 22, after which they are converted into sinks. Before layer 22, their hidden-state, key, and value norms follow the average-token trend; afterward, they closely track the BOS sink, whose hidden-state norm is several orders of magnitude higher than that of normal tokens, while its key and value norms are lower.
Not all models exhibit secondary sinks. For example, although multiple sinks also appear in LLaMA-7B-Chat Figure˜1 (Bottom), they are exclusively primary sinks. Among the 11 model families we analyzed, secondary sinks were found in Qwen2, Qwen2-Math, Qwen2.5, Qwen2.5-Math, Qwen3, and QwQ. Full detection results on all 11 model families can be found in Appendix A. These sinks typically occur in semantically uninformative tokens Table˜3 and can appear at any position during the generation Figure˜5.
@@ 4 Causal Formation Of Secondary Sinks
Figure˜3 (Right) shows the cosine similarity of 75 future secondary sink tokens from AIME24 in DeepSeek-14B (layer 22). The MLP inputs ${\bm{x}}_{t}$ initially have low similarity to the BOS direction, but this similarity increases through the MLP; by the output ${\bm{f}}_{t}$, the vectors are nearly fully aligned with the BOS sink.
This suggests that the sink direction is encoded in the MLP and that future sink tokens contain enough information to recover it. Following the linear hypothesis (Park et al., 2023), we hypothesize that these tokens share a common key that retrieves the sink direction. We test this by performing PCA on the concatenated MLP inputs ${\bm{X}}^{l}$ from 250 sink tokens. The spectrum is strongly low-rank. Feeding the top individual principal components through the MLP shows that these PCs are mapped to the same sink direction. Figure˜7 plots the MLP output norm and cosine similarity with the BOS sink when feeding $\pm\alpha\mathrm{PC}_{i}$. We find that the MLP amplifies components aligned with the BOS sink and suppresses those that are misaligned, producing the large hidden-state norms.
@@
To better isolate the decision-making contributions from attention, MLP, and the hidden state, we substitute their outputs at future sink tokens with the average semantically uninformative token. See Appendix C for further details.
@@ 5 Secondary Sinks Effect on Attention
As shown in  Figure˜4 (Left) , the sink-score for the BOS sink shows a valley shape across the model depth, and it is when it reaches its lowest sink score that the secondary sinks started emerging. This suggests the secondary sinks may serve to compensate for the decay in the BOS sink.
@@
We observe that secondary sinks persist for varying depths, ranging from as few as 2 layers to as many as 22 layers, approximately half of the network. As shown in Figure˜4 (right), both the sink lifetime and the sink score of the secondary sink are strongly correlated with the $\ell_{2}$-norm of the MLP output ${\bm{f}}^{22}_{t}$ at $l_{\text{start}}$. The ratio between the sink scores of the secondary sink and the BOS sink exhibits a log-linear relationship with ${\bm{f}}^{22}_{t}$, with the largest slope observed at layer 33. In contrast, the sink lifetime increases monotonically with the log-norm, exhibiting distinct linear regimes separated by plateaus.
@@ Appendix A Additional Detection Results
Appendix B Properties of Secondary Sinks
Figure 5: Distribution of sinks across token index.
Model Secondary Sinks Token Distribution
Deepseek-14b " " "1" "^ " "2" "," "(" " I"
56.41% 17.95% 5.13% 5.13% 2.56% 2.56% 2.56%
[…]
19.87% 19.87% 13.91% 9.93% 4.64% 3.97% 3.97%
[…]
6.13% 5.28% 3.62% 3.56% 3.50% 3.48% 2.98%
Qwen2.5-14B " " "0" "1" "8" "=" "_" "t"
74.71% 9.20% 5.75% 2.30% 2.30% 1.15% 1.15%
Table 3: Frequency of secondary sink tokens.
Appendix C Causal formation of Secondary Sinks
Token Swapping